### v2

# VITA: Vision-to-Action Flow Matching Policy

Abstract
Conventional flow matching and diffusion-based policies sample through iterative denoising from standard noise distributions (e.g., Gaussian), and require conditioning mechanisms to incorporate visual information during the generative process, incurring substantial time and memory overhead. To reduce the complexity, we develop VITA (VIsion-To-Action policy), a noise-free and conditioning-free policy learning framework that directly maps visual representations to latent actions using flow matching. VITA treats latent visual representations as the source of the flow, thus eliminating the need of conditioning. As expected, bridging vision and action is challenging, because actions are lower-dimensional, less structured, and sparser than visual representations; moreover, flow matching requires the source and target to have the same dimensionality. To overcome this, we introduce an action autoencoder that maps raw actions into a structured latent space aligned with visual latents, trained jointly with flow matching. To further prevent latent space collapse, we propose flow latent decoding, which anchors the latent generation process by backpropagating the action reconstruction loss through the flow matching ODE (ordinary differential equations) solving steps. We evaluate VITA on 8 simulation and 2 real-world tasks from ALOHA and Robomimic. VITA outperforms or matches state-of-the-art generative policies, while achieving $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional methods with conditioning. Project page: VITA.
Keywords Imitation Learning $\cdot$ Flow Matching $\cdot$ Robotics Manipulation $\cdot$ Policy Learning
1 Introduction
Figure 1: A comparison between VITA and conventional flow matching and diffusion policies. Unlike conventional methods that sample noise from standard distributions and inject source input modalities via conditioning during generation, VITA poses no constraints on the source distributions, and flow directly from latent visual representations to latent actions, eliminating the need for conditioning.
Flow matching and diffusion models have demonstrated remarkable success across a wide range of cross-modal generation tasks, from text-to-image generation stablediffusion ; dit ; sit ; crossflow ; flowtok ; conddiffusion , text-to-video generation video-diff ; li2023videogen ; jin2024pyramidal , to visuomotor (vision-to-action) policy chi2023diffusion ; ren2024diffusion ; gao2025ril ; su2025dense ; fmp-3d-consist ; fmp-multisupport ; fmp-riemannian ; black2410pi0 . Conventional flow matching and diffusion methods  lipman2023flowmatchinggenerativemodeling ; sohl2015deep generate samples by starting with noise sampled from a basic source distribution (often Gaussian) and progressively “denoising” them into the target modality, requiring additional conditioning mechanisms  stablediffusion ; conddiffusion ; chi2023diffusion ; dit-policy to inject visual modality at denoising steps, which incurs increased time and space complexity crossflow ; flowtok .
Clearly, minimizing complexity is critical for effective real-time robot control. To overcome the inefficiencies inherent to conditioning mechanisms in conventional methods, we develop VITA (VIsion-To-Action flow matching), a noise-free policy learning framework that directly maps visual representations to latent actions. As depicted in  Figure˜1, unlike conventional methods that flow from a Gaussian prior and inject visual contexts via conditioning mechanisms during generation, VITA poses no constraints on the source distribution, and directly flows from latent visual representations, obviating the need for injecting visual inputs via conditioning modules. Consequently, VITA significantly reduces inference latency and memory overhead, simplifying the network architecture.
Learning vision-to-action flow matching, however, presents several new challenges. In general, bridging two distinct modalities is inherently difficult crossflow ; flowtok , particularly in robot learning where action data is relatively limited, unstructured and sparse, whereas the visual representations exhibit rich structures and semantics and have far higher dimensions. Additionally, flow matching requires that the source and target have equal dimensionalities, which prevents directly pairing raw actions with visual representations.
To address these challenges, we propose two key designs for VITA. 1) Learned target latent actions for flow matching. We introduce a structured latent action space, learned via an action autoencoder, that ‘lifts’ action representations to match the higher dimensionality of visual representations and serves as a structured target distribution for flow matching. The action encoder up-samples raw actions into target latent actions, and a decoder reconstructs raw actions from these latents. 2) Flow latent decoding. In conventional flow matching methods such as latent diffusion for image generation (stablediffusion, ), the target latent space can be pre-trained with abundant image data and then frozen as reliable flow matching targets; in contrast, we show that a pre-trained and frozen latent action space for learning vision-to-action flows yields poor performance, since action data can be too sparse and limited to learn reliable target latents and frozen latent spaces cannot be corrected. It is therefore plausible to jointly train the flow model with the action autoencoder. To prevent model collapse of target latent action space (which happens by naively reducing the flow matching and autoencoder loss), we introduce flow latent decoding which backpropagates the reconstruction loss of latent actions generated by solving flow matching ordinary differential equations (ODE), anchoring the latent generation process using ground-truth actions. This approach also bridges the training-inference gap of latent actions: during training, the flow matching model learns to match the targets given by the action encoder, and the decoder is trained to reconstruct these targets. whereas at test time, the decoder must reconstruct ODE-generated latent actions.
We evaluate VITA on both real-world and simulated tasks using ALOHA chuang2024active ; fu2024mobile ; zhao2024aloha and Robomimic robomimic2021 . Our experiments demonstrate that VITA outperforms or matches state-of-the-art generative policies. We implement VITA with only MLP layers. Compared to state-of-the-art flow matching and diffusion policies that predominantly leverage complex architecture such as transformers sit or U-Net unet , and introduce conditioning modules zhao2023learning ; fmp-3d-consist , the MLP-only and noise-free VITA policy achieves, $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional flow matching. To our knowledge, VITA is the first MLP-only flow matching policy to succeed on tasks as challenging as ALOHA bi-manual manipulation.
Our main contributions are summarized as follows:
Noise-Free Flow Matching for Visuomotor Learning. We propose VITA, a noise-free policy that directly evolves latent visual representations into latent actions via flow matching. To bridge the modality gap, VITA incorporates an action autoencoder that maps raw actions into structured latent spaces aligned with visual representations. To prevent latent space collapse, we propose flow latent decoding, which anchors latent generation using ground-truth actions by backpropagating the action reconstruction loss through the flow matching ODE (ordinary differential equations) solving steps.
An Efficient Policy Architecture. By treating vision as the source distribution, VITA obviates costly conditioning mechanisms required by conventional diffusion and flow matching policies. VITA also enables a lightweight MLP-only implementation that achieves superior performance on challenging bi-manual manipulation tasks with high-dimensional visual inputs.
State-of-the-Art Performance in Real and Simulated Environments. We validate VITA’s effectiveness on 8 simulation and 2 real-world tasks. VITA surpasses or matches state-of-the-art generative policies in success rates while delivering $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional flow matching approaches.
2 Related Work
Imitation Learning for Visuomotor Policy. Imitation learning enables robots to learn complex behaviors by mimicking expert demonstrations. Behavior cloning is a prominent imitation learning paradigm that learns a policy that maps observations to actions (zhao2023learning, ; lee2024interact, ) via supervised learning. Recent advancements in behavioral cloning have widely adopted autoregressive modeling (fu2024context, ; carp, ; su2025dense, ) and generative modeling (chi2023diffusion, ). Generative modeling learns a conditional distribution of actions given an observation, leveraging conditional variational autoencoders (CVAEs) (zhao2023learning, ; lee2024interact, ), diffusion (dit-policy, ; chi2023diffusion, ), or flow matching (fmp-affordance, ; fmp-3d-consist, ). Generative models ubiquitously require conditioning modules (e.g., cross-attention (dit-policy, ), AdaLN (dit-policy, ), FiLM (perez2018film, ; chi2023diffusion, )) to inject observations at each step of the generation process. VITA removes the visual conditioning module by developing a noise-free and conditioning-free vision-to-action flow.
Diffusion and Flow matching for Generative Modeling. Diffusion and flow matching models have shown strong performance across diverse generative tasks, such as image and video generation (stablediffusion, ; video-diff, ), and visuomotor policies (chi2023diffusion, ). Unlike diffusion which samples from a Gaussian distributions, flow matching theoretically places no constraints on the choice of source distribution (tong2024bridgesscore, ). A few works have explored this property to learn the direct transport within the same modality (i2i-flow-interpolants, ; i2i-flow-sim-free, ). Recently,  crossflow and  flowtok extended this to more challenging cross-modal generation between text and image. VITA focuses on learning to bridge vision and action for visuomotor control, where the action modality presents unique challenges because of limited data and its unstructured nature. Different from flow matching for image generation, which typically pre-trains and freezes the latent image space when learning the flow (stablediffusion, ; flowtok, ; crossflow, ), VITA resorts to a fully end-to-end pipeline training to effectively learn the latent action space from limited and sparse action data along with flow matching. We propose flow latent decoding to backpropagate action reconstruction losses through the the latent action generation process (ODE solving steps) during training.
3 Preliminaries
Flow matching models learn to transport samples from a source probability distribution $p_{0}$ to a target distribution $p_{1}$ by learning a velocity vector field $v_{\theta}$ (lipman2023flowmatchinggenerativemodeling, ; rectified-flow, ). This generative process is defined by an ordinary differential equation (ODE), $\frac{d\bm{z}_{t}}{dt}=v_{\theta}(\bm{z}_{t},t)$, where $t\in[0,1]$ is a continuous time variable, $\bm{z}_{t}$ represents a sample at time $t$, and the model $v_{\theta}$ is a neural network. The goal of training is for $v_{\theta}$ to generate a path that transforms a sample $\bm{z}_{0}\sim p_{0}$ into a corresponding sample $\bm{z}_{1}\sim p_{1}$.
For a simple, straight probability path between two samples, the interpolated state at time $t$ is defined as $\bm{z}_{t}=(1-t)\bm{z}_{0}+t\bm{z}_{1}$ The corresponding ground-truth velocity of this path given by $u_{t}=\frac{d\bm{z}_{t}}{dt}=\bm{z}_{1}-\bm{z}_{0}$. The network $v_{\theta}$ is trained by minimizing the Mean Squared Error (MSE) between its predicted velocity and the ground-truth velocity:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{t,\bm{z}_{0},\bm{z}_{1}}\left[\left\|v_{\theta}(\bm{z}_{t},t)-(\bm{z}_{1}-\bm{z}_{0})\right\|^{2}\right]$ (1)
During inference, starting from a source sample $\bm{z}_{0}$, the model generates a target sample $\hat{\bm{z}}_{1}$ by solving the ODE initial value problem from $t=0$ to $t=1$:
$\hat{\bm{z}}_{1}=\bm{z}_{0}+\int_{0}^{1}v_{\theta}(\bm{z}_{t},t)\,dt$ (2)
Notably, conventional flow matching or diffusion policies for visuomotor control generate actions by evolving a sample from a simple prior distribution, such as Gaussian noise ($\bm{z}_{0}\sim\mathcal{N}(0,I)$), and require conditioning on visual inputs to guide the generation process (fmp-affordance, ; chi2023diffusion, ; dit-policy, ). However, flow matching does not assume the source distribution must be simple noise; it can theoretically connect two arbitrary distributions. We exploit this flexibility to re-formulate generative visuomotor policy learning: we directly use the latent distribution of visual observations as the source distribution $p_{0}$, yielding a noise-free flow matching process that eliminates the need for costly conditioning modules and therefore boosts time and space efficiency.
4 VITA: Vision-to-Action Flow Matching
The key challenge in VITA is the large dimensionality gap between vision and action, compounded by the sparsity and unstructured nature of action data. In this section, we present the core designs of VITA developed to address these issues. We first introduce the mathematical formulation of VITA (Section˜4.1) and its overall architecture (Section˜4.2). We then show why constructing a latent action space is essential for resolving dimensionality mismatch (Section˜4.3), and propose flow latent decoding to address model collapse. Finally, we describe the objectives that enable effective end-to-end VITA learning from scratch (Section˜4.4).
4.1 Flowing from Vision to Action
VITA learns a policy $\pi(A|O)$ that directly maps observations $O$ to a corresponding sequence of future actions $A$. The observations $O$ encompasses raw visual inputs $I\in\mathbb{R}^{H\times W\times C}$ and, optionally, the robot’s proprioceptive state vector $S$. Actions are represented as temporal sequences over a prediction horizon, formally defined as $A\in\mathbb{R}^{T_{\text{pred}}\times D_{\text{action}}}$, where $T_{\text{pred}}$ is the prediction horizon and $D_{\text{action}}$ is the dimensionality of the action space. We employ action chunking with $T_{\text{pred}}>1$ to enhance temporal consistency  (zhao2023learning, ).
Unlike conventional flow matching policies that generate actions through conditional denoising of random noise given visual inputs, VITA formulates the visuomotor policy as an unconditional flow between the source distribution of latent visual representations $p_{0}$ and the target distribution of latent actions $p_{1}$. The flow matching framework learns to evolve a latent visual representation $\bm{z}_{0}$ into an latent action representation $\bm{z}_{1}$ by learning a velocity field $v_{\theta}(\bm{z}_{t},t)$, where $t\in[0,1]$ represents the continuous time parameter interpolating between the source state at $t=0$ and the target state at $t=1$. Critically, flow matching requires $\bm{z}_{0}$ and $\bm{z}_{1}$ to share identical dimensionality, necessitating the construction of a latent action space that matches the dimensionality of visual representations (Section˜4.3).
During inference, the current observation $O_{\text{curr}}$ is first encoded into its latent visual representation $\bm{z}_{0}=\mathcal{E}_{v}(O_{\text{curr}})$. This latent vector is subsequently evolved into a predicted action latent representation, $\hat{\bm{z}}_{1}$, by numerically solving the ODE from $t=0$ to $t=1$ using the learned velocity field $v_{\theta}$: $\hat{\bm{z}}_{1}=\bm{z}_{0}+\int_{0}^{1}v_{\theta}(\bm{z}_{t},t)\,dt$. In other words, $\hat{\bm{z}}_{1}$ is obtained by numerically solving the flow matching ODE, and serves as an approximation to the target latent $\bm{z}_{1}$. The resulting latent action vector $\hat{\bm{z}}_{1}$ is then decoded through the action decoder to yield the final action sequence $\hat{A}=\mathcal{D}_{a}(\hat{\bm{z}}_{1})$.
4.2 VITA Architecture Design
Figure 2: An overview of the VITA architecture: The vision encoder maps observations into a source latent representation $\bm{z}_{0}$ for the flow; the action encoder provides a target latent representation $\bm{z}_{1}$ for flow matching training. The action decoder learns to decode $\hat{\bm{z}}_{1}$ (latent actions generated by solving ODEs) to actions via flow latent decoding losses, and decode $\bm{z}_{1}$ to actions (latent actions from action encoder) via AE losses. The flow matching network is designed to learn the velocity field over a continuous flow matching path from $\bm{z}_{0}$ to $\bm{\hat{z}}_{1}$.
As depicted in Figure˜2, VITA  is composed of the three primary components: 1) The Vision Encoder ($\mathcal{E}_{v}$) maps raw camera images into a vector representation of visual latents $I$. The observation $O$ to the policy consists of both image vector representations $I$ and, optionally, the robot’s proprioceptive states $S$. In short, $O$ is mapped into a latent vector $\bm{z}_{0}=\mathcal{E}_{v}(O),\quad\text{where }\bm{z}_{0}\in\mathbb{R}^{D_{\text{latent}}}$ is the source of flow matching. 2) The Action Autoencoder (AE) consists of the Action Encoder and the Action Decoder, and learns a compact representation for action chunks. The Action Encoder ($\mathcal{E}_{a}$) maps the ground-truth action chunk $A$ to latent actions $\bm{z}_{1}=\mathcal{E}_{a}(A),\quad\text{where }\bm{z}_{1}\in\mathbb{R}^{D_{\text{latent}}}$ serves as the target for flow matching; the Action Decoder ($\mathcal{D}_{a}$) reconstructs an action chunk $\hat{A}=\mathcal{D}_{a}(\hat{\bm{z}}_{1})$ from latent actions $\hat{\bm{z}}_{1}$. 3) The Flow Matching Network ($v_{\theta}$) is learned to predict the velocity field at arbitrary $t$.
4.3 Bridging the Modality Gap between Vision and Action
A key constraint of flow matching is that the source and target distributions must share the same dimensionality. This poses a critical challenge for vision-to-action policies, since action spaces are typically much lower-dimensional than visual representations. For example, action dimensionalities range from 2 on PushT to 21 on ThreadNeedle, whereas visual representations can be 512-dimensional (zhao2024aloha, ) or even higher when using grid-based tokenizers (chi2023diffusion, ).
To bridge the gap, one naive option is to down-sample latent visual representations to action chunk dimensionalities which, however, causes severe information loss and degrades performance. Alternatively, one can up-sample actions with zero-padding, yielding sparse, unstructured targets that hinder flow matching learning (see Appendix B.1). A third alternative is a pre-trained, frozen action AE, akin to common practice in latent diffusion for image generation (stablediffusion, ), but this proves ineffective for learning vision-to-action flow: with sparse, limited action data the induced latent space is unreliable as a flow target and cannot be corrected once frozen. As another alternative, jointly training the action AE with flow matching may still fail, and our empirical studies identify the root cause as latent space collapse induced by a training-inference gap in the latent actions used for decoding as detailed below.
Training-Inference Gap between Encoder-Based and ODE-Generated Latent Actions.
During training, the decoder reconstructs actions from encoder-based latent actions $\bm{z}_{1}$, whereas at inference it decodes $\hat{\bm{z}}_{1}$ generated by solving the flow matching ODE. Since $\hat{\bm{z}}_{1}$ is an approximation, and does not always align with $\bm{z}_{1}$, the decoder can fail to map them into meaningful actions. To address this gap, we propose flow latent decoding, which enforces the model to decode from ODE-generated latent actions $\hat{\bm{z}}_{1}$ during training, anchoring the latent generation process with ground-truth actions.
4.4 VITA Learning Objectives
Building upon our analysis of the training-inference gap, we now formulate a comprehensive learning framework for VITA that prevents latent collapse while ensuring effective end-to-end optimization. Our framework includes three essential objectives: flow latent decoding (FLD), flow matching (FM), and action autoencoder (AE) losses, each addressing distinct aspects of the learning challenge.
Flow Latent Decoding (FLD).
FLD addresses the training-inference gap by anchoring ODE-generated actions using ground-truth actions during training. Formally, FLD is defined as the reconstruction loss using ODE-generated latent actions, $\mathcal{L}_{\mathrm{FLD}}=\big\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\big\|$, where $\hat{\bm{z}}_{1}$ is obtained by solving the flow ODE with an Euler solver during training. FLD propagates gradients through the decoder and the ODE solver into both $v_{\theta}$ and $\mathcal{E}_{v}$. By decoding $\hat{\bm{z}}_{1}$ into actions and measuring reconstruction error directly in action space, FLD effectively minimizes the discrepancies between encoder-based and ODE-generated latents.
Flow Latent Consistency (FLC).
To gain deeper insight into the mechanics of FLD, we introduce flow latent consistency (FLC), a minimalist surrogate that directly aligns ODE-generated and encoder-based latents without decoding. Formally, FLC is defined as $\mathcal{L}_{\mathrm{FLC}}=\big\|\hat{\bm{z}}_{1}-\bm{z}_{1}\big\|$. Under mild local regularity assumptions on $\mathcal{D}_{a}$ (stated below), FLC and FLD provide locally equivalent training signals for the same $\hat{\bm{z}}_{1}$. Empirically, FLC also prevents collapse without explicit action reconstruction, though convergence is slightly slower than with FLD (Section˜5.3). This theoretical connection not only validates our approach but also offers computational flexibility in implementation. A sketch of the analysis is given below, with full proofs and corollaries deferred to Appendix A.
Assumption 1 (Decoder locally well-behaved).
$\mathcal{D}_{a}$ is $C^{1}$ in a neighborhood of $\bm{z}_{1}$, with Jacobian singular values bounded as $m\!\leq\!\sigma_{\min}\!\leq\!\sigma_{\max}\!\leq\!L$ on that neighborhood. Let $\varepsilon_{\mathrm{AE}}:=\|\mathcal{D}_{a}(\bm{z}_{1})-A\|$ denote the local AE reconstruction error.
Theorem 1 (Local equivalence of FLD and FLC).
Under ˜1, for any $\hat{\bm{z}}_{1}$ in the neighborhood,
$m\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|-\varepsilon_{\mathrm{AE}}\ \leq\ \|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|\ \leq\ L\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|+\varepsilon_{\mathrm{AE}}.$
If $\varepsilon_{\mathrm{AE}}=0$, the minimizers of $\mathcal{L}_{\mathrm{FLD}}$ and $\mathcal{L}_{\mathrm{FLC}}$ coincide and equal $\{\bm{z}_{1}\}$; if $\varepsilon_{\mathrm{AE}}>0$, any minimizer of $\mathcal{L}_{\mathrm{FLD}}$ lies within radius $\varepsilon_{\mathrm{AE}}/m$ of $\bm{z}_{1}$.
This theoretical result confirms that FLD and FLC target the same underlying optimization objective.
We will provide further discussion in Section˜5.3, showing that including the flow latent decoding loss with a non-zero $\lambda_{\text{FLD}}$ is critical for avoiding latent space collapse and training successful VITA policies. We will also ablate the effects of $\lambda_{\text{FLD}}$, $\lambda_{\text{FLC}}$, and $\lambda_{\text{AE}}$ in Figure˜6.
Flow Matching and Autoencoder Losses.
The flow matching loss supervises the flow network $v_{\theta}$ by minimizing the MSE between the predicted velocity and the ground-truth velocity as shown in Equation˜1. The action autoencoder loss trains $(\mathcal{E}_{a},\mathcal{D}_{a})$ to reconstruct action chunks using an L1 loss, $\mathcal{L}_{\text{AE}}=\|A-\mathcal{D}_{a}(\mathcal{E}_{a}(A))\|_{1}$, where $\bm{z}_{1}=\mathcal{E}_{a}(A)$ serves as a structured target latent with small reconstruction bias and good local conditioning. This structured latent space strengthens the theoretical link between FLD and FLC and further stabilizes training.
The training objective can be written as a weighted sum of all three losses: $\mathcal{L}_{\text{VITA}}=\lambda_{\text{FM}}\mathcal{L}_{\text{FM}}+\lambda_{\text{FLD}}\mathcal{L}_{\text{FLD}}+\lambda_{\text{AE}}\mathcal{L}_{\text{AE}}$.
5 Experiments
We evaluate VITA on ALOHA (chuang2024active, ; fu2024mobile, ) across 5 simulation tasks (Figure˜3) and 2 real-world tasks (Figure˜4). The simulation environments and datasets are from AV-ALOHA (chuang2024active, ), which extend ALOHA with an additional 7-DoF arm equipped with an active-vision camera that dynamically adjusts its viewpoint. This setup increases task difficulty due to the larger action space and non-stationary observation dynamics. In addition, we evaluate on 2 single-arm Robomimic simulation tasks.
For simulation tasks, each environment provides 100-200 demonstrations. Demonstrations were collected from human experts in simulation using VR headsets, with the left-eye image used as input. VITA was trained at $25/3$ FPS on AV-ALOHA and $20$ FPS on Robomimic. For real-world tasks on AV-ALOHA, (Figure˜4). Each task is trained with 50 demonstrations, using the left image from the stereo active-vision camera as input. VITA was trained at 11 FPS, and linearly interpolated to 33 FPS for deployment on the physical robot.
Figure 3: An illustration of 5 AV-ALOHA tasks (CubeTransfer, SlotInsertion, HookPackage, PourTestTube, ThreadNeedle), and 2 Robomimic tasks (Square, Can).
Figure 4: An illustration of two challenging real-world AV-ALOHA tasks, HiddenPick, and TransferFromBox. The pictures are taken from autonomous rollouts by the VITA policy.
5.1 Implementation
Flow Matcher. We adopt OT-CFM (ot-cfm, ), a conditional flow matching method based on optimal transport, and solve the Euler ODE with 6 steps using linearly interpolated $t$ for both VITA and FM baselines.
Networks. We use the widely adopted ResNet-18 (resnet, ; su2025dense, ; chi2023diffusion, ) as the vision encoder for all the experiments. We use the 512-dimensional feature vector after the average pooling layer as the visual representations for VITA. Therefore, the source $\bm{z}_{0}$ and target latents $\bm{z}_{1}$ are both 512-dimensional vectors ($D_{\text{latent}}=512$). The vector representations and conditioning-free design is amenable to simple and efficient implementation. For example, we use 4-layer MLP-only networks for both flow matching and decoding.
Training & Evaluation. We train VITA on each task for 50,000 steps with a batch size of 128. The policy predicts action chunks of length 16, of which the first 8 actions are executed. In simulation, we evaluate the policy every 500 steps, with each evaluation consisting of 50 episodes. We report the best success rates (SRs), averaged over 3 random seeds, for both VITA and baselines (Table˜1). For real-world tasks, we evaluate 3 checkpoints over 20 episodes each and report the best SRs (Table˜2).
Baselines. We compare VITA against state-of-the-art generative policies, including flow matching (FM) policy (fmp-affordance, ), diffusion policy (DP)(chi2023diffusion, ), and action chunking transformer (ACT)(zhao2023learning, ). We follow the ACT and DP implementation in LeRobot cadene2024lerobot . We found DP requires significantly more training steps than FM; ACT yields sub-optimal performance on most tasks; therefore, we increase training steps to 200,000 for both DP and ACT, and then report the best success rates. For DP, each inference takes 10 DDPM (Denoising Diffusion Probabilistic Model) steps (chi2023diffusion, ). We evaluate both the task performance and inference efficiency of VITA relative to these baselines in Section˜5.2.
5.2 Performance
Success Rates.
We evaluate VITA against state-of-the-art conditional generative policies on 8 simulated tasks: 5 bimanual AV-ALOHA tasks with active vision (chuang2024active, ), 2 single-arm Robomimic tasks (robomimic2021, ), and 1 PushT task. Results are reported in Table˜1. Across all simulation tasks, VITA consistently outperforms or matches state-of-the-art baselines. ACT shows limited performance even after convergence, while DP requires significantly more training steps and remains sub-optimal despite extended training. We further assess VITA in real-world settings on two challenging AV-ALOHA tasks, with performance summarized in Table˜2 and examples of autonomous rollouts in Figure˜4, highlighting the precision of vision-to-action flow in manipulation tasks with high-dimensional visual inputs and large action spaces.
Table 1: Task success rates (SR) across simulation tasks including ALOHA, Robomimic, and PushT. We report the best SR during validation, and compute the mean across 3 random seeds.
Task VITA FM DP ACT
ThreadNeedle 0.92 0.88 0.00 0.52
SlotInsertion 0.76 0.83 0.20 0.50
PourTestTube 0.79 0.84 0.06 0.14
HookPackage 0.86 0.83 0.12 0.22
CubeTransfer 1.00 1.00 0.76 0.98
PushT 0.88 0.84 0.80 0.30
Square 1.00 1.00 1.00 0.46
Can 1.00 1.00 1.00 0.78
Table 2: Success rates of VITA on two real-world AV-ALOHA tasks. Each task has 3 subtasks.
HiddenPick TransferFromBox
Subtask Reveal Pick Place Pick Transfer Place
SR 1.00 0.65 0.65 1.00 0.95 0.90
Efficiency.
The VITA policy is highly efficient due to its architectural simplicity. The core components used during inference, the flow matching network and the action decoder, are both lightweight MLP-only networks. We use vector representations for both vision and action, further reducing time and space overhead. VITA contains 31M parameters, including the ResNet-18. We benchmark the inference latency and policy throughput of VITA and compare it against conventional flow matching policy baselines. We implement two different flow matching parameterizations, including diffusion transformer (dit, ), and U-Net (unet, ), and three different conditioning mechanisms, including AdaLN (dit, ; dit-policy, ), cross-attention (attention, ; dit-policy, ; carp, ), and FiLM (perez2018film, ; chi2023diffusion, ), as the baselines. FM and DP are typically parameterized using U-Nets (unet, ) or diffusion transformers (DiTs) (dit, ), and are trained to predict velocity fields or noise at each denoising step. While U-Nets and DiTs are highly expressive, they can be large or computationally expensive, posing a drawback for real-world robotic deployments that require low-latency inference. Most importantly, these methods inevitably incorporate visual conditioning mechanisms, such as cross, further increasing both inference time and memory footprint.
With a control frequency of $25/3\approx 8.33$ Hz, our MLP-only VITA policy achieves an inference wall-clock time of 0.22 ms on an NVIDIA 4090 GPU. In contrast, a comparable transformer-based flow matching policy incurs higher latency (0.33-0.51 ms) and requires more parameters. A detailed breakdown of efficiency metrics is given in Table˜3.
Table 3: Model sizes and inference latency of different policy implementations.
Model Conditioning Architecture Num of Params Latency (ms/chunk)
VITA N/A MLP 31.09M 0.2215
FM AdaLN DiT 31.16M 0.3307
FM Cross-Attn DiT 29.06M 0.5102
FM FiLM U-Net 84.05M 0.3650
DDPM FiLM U-Net 81.82M 2.5985
5.3 Ablation of Flow Latent Decoding
We investigate the importance of flow latent decoding ($\mathcal{L}_{\text{FLD}}$) for effectively training VITA policies. During training, the decoder is trained by reconstructing from encoder-based latent actions ${\bm{z}}_{1}$, while during inference, the decoder needs to reconstruct from ODE-generated latent actions $\hat{\bm{z}}_{1}$. As a result, the generated latent actions at inference may not decode into meaningful actions, as demonstrated by  Figure˜6. To bridge this gap, we propose two objectives to minimize the discrepancy between encoder-based latents ${\bm{z}}_{1}$ and ODE-generated latents $\hat{\bm{z}}_{1}$: (1) FLD, which decodes $\hat{\bm{z}}_{1}$ into the action space and computes the reconstruction error against the ground-truth action; and (2) FLC, which directly aligns the predicted latent action with the encoded latent target $\bm{z}_{1}=\mathcal{E}_{a}(A)$ using $\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|$.
Figure 5: Comparison of reconstructed actions between (a) VITA, and (b) VITA without FLD. Reconstruction fails without FLD because of latent space collapse.
Figure 6: Success rates using different objectives.
As shown in Figure˜6, without flow latent decoding (i.e., $\lambda_{\text{FLD}}=0$), the model completely fails to learn a reasonable policy. In contrast, applying FLD succeeds in learning the policy. FLC provides a weaker signal compared to FLD (because FLD directly anchors using the ground-truth actions), resulting in slightly slower convergence. However, we found that using a combination of both objectives yields the best performance due to richer learning signals on both action space and latent space. In short, FLD and FLC are crucial for learning VITA, and in practice a combination of both can achieve optimal results.
5.4 VITA Denoising Process
Figure 7: Conventional flow matching (a) flows from random Gaussian noise, and gradually “denoise" into raw action chunks by solving ODEs; in contrast, VITA (b) flows from latent images to latent actions, and the latent actions are then decoded to action chunks.
Figure˜7 compares the denoising processes of VITA and conventional flow matching. By jointly optimizing the vision-to-action flow and an action autoencoder with FLD, VITA induces highly structured latent spaces for both vision and action representations. As a result, even the latent image can be directly decoded into a coherent trajectory using the action decoder, which is then progressively refined by the VITA flow matching ODE.
6 Conclusion
In this work, we developed VITA, an efficient and high-performing visuomotor policy. VITA generates actions in a noise-free manner and directly evolves latent visual representations into latent actions. By leveraging latent image representations as the flow’s source distribution, VITA eliminates the need for complex conditioning mechanisms such as cross-attention, significantly simplifying model architecture and enhancing efficiency. VITA employs vector representations for both latent images and actions, thereby enabling a simple MLP-only architecture for vision-to-action flow. Two key designs are critical to VITA’s success — structuring the latent action space using an action autoencoder and enabling backpropagation of flow latent decoding losses across ODE solving steps — allowing VITA to effectively learn latent action spaces with flow matching in a fully end-to-end manner. Extensive experiments demonstrate that VITA achieves state-of-the-art performance on ALOHA and Robomimic in both real and simulation tasks. VITA achieves $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional flow matching using different conditioning mechanisms.
Appendix A Proof of FLD and FLC Equivalence
Preliminaries.
All norms below are vector norms with induced operator norms. We use the ball $B(\bm{z}_{1},r)=\{\bm{x}:\|\bm{x}-\bm{z}_{1}\|<r\}$. Assumption 1 in the main text holds throughout.
Lemma 1 (Local bi-Lipschitzness from Jacobian bounds).
For any $\hat{\bm{z}}_{1}\in B(\bm{z}_{1},r)$,
$m\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|\ \leq\ \|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-\mathcal{D}_{a}(\bm{z}_{1})\|\ \leq\ L\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|.$
Proof of Lemma 1. Let $\gamma(s)=\bm{z}_{1}+s(\hat{\bm{z}}_{1}-\bm{z}_{1})$ for $s\in[0,1]$. The mean-value integral formula gives
$\mathcal{D}_{a}(\hat{\bm{z}}_{1})-\mathcal{D}_{a}(\bm{z}_{1})=\int_{0}^{1}J_{\mathcal{D}_{a}}(\gamma(s))\,(\hat{\bm{z}}_{1}-\bm{z}_{1})\,ds.$
Taking norms and using, for any matrix $J$ and vector $v\neq 0$, $\sigma_{\min}(J)\|v\|\leq\|Jv\|\leq\sigma_{\max}(J)\|v\|$, together with $m\leq\sigma_{\min}(J_{\mathcal{D}_{a}}(\gamma(s)))$ and $\sigma_{\max}(J_{\mathcal{D}_{a}}(\gamma(s)))\leq L$ for all $s$, yields the bounds.
Proof of Theorem 1. Add and subtract $\mathcal{D}_{a}(\bm{z}_{1})$ and apply triangle/reverse-triangle inequalities:
$\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|\leq\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-\mathcal{D}_{a}(\bm{z}_{1})\|+\|\mathcal{D}_{a}(\bm{z}_{1})-A\|,$
$\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|\geq\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-\mathcal{D}_{a}(\bm{z}_{1})\|-\|\mathcal{D}_{a}(\bm{z}_{1})-A\|.$
Invoke Lemma 1 and set $\varepsilon_{\mathrm{AE}}=\|\mathcal{D}_{a}(\bm{z}_{1})-A\|$ to obtain the two-sided inequality stated in Theorem 1. The minimizer claims follow immediately: if $\varepsilon_{\mathrm{AE}}=0$, both losses are minimized at $\hat{\bm{z}}_{1}=\bm{z}_{1}$; otherwise any minimizer of FLD must satisfy $\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|\leq\varepsilon_{\mathrm{AE}}/m$.
Corollary A.1 (squared-loss version).
Assume $\varepsilon_{\mathrm{AE}}=0$. Then
$m^{2}\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|^{2}\ \leq\ \|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|^{2}\ \leq\ L^{2}\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|^{2}.$
With $\varepsilon_{\mathrm{AE}}=0$, Lemma 1 gives $m\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|\leq\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-\mathcal{D}_{a}(\bm{z}_{1})\|\leq L\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|$. Since both sides are nonnegative, squaring preserves the inequalities. Thus the squared FLD objective is sandwiched between $m^{2}$ and $L^{2}$ times the squared FLC objective. Consequently, the map $\hat{\bm{z}}_{1}\mapsto\|\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\|_{2}^{2}$ is locally $L^{2}$-smooth and $m^{2}$-strongly convex along latent directions (intuitively, its curvature is controlled by $J_{\mathcal{D}_{a}}^{\top}J_{\mathcal{D}_{a}}$ whose eigenvalues lie in $m^{2},L^{2}]$).
Corollary A.2 (gradient scaling for squared losses).
Let $J:=J_{\mathcal{D}_{a}}(\hat{\bm{z}}_{1})$ and assume $\varepsilon_{\mathrm{AE}}=0$. For the squared losses,
$\nabla_{\hat{\bm{z}}_{1}}\mathcal{L}^{(2)}_{\mathrm{FLD}}=2\,J^{\top}\!\big(\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A\big),\qquad\nabla_{\hat{\bm{z}}_{1}}\mathcal{L}^{(2)}_{\mathrm{FLC}}=2\,(\hat{\bm{z}}_{1}-\bm{z}_{1}).$
Then
$m^{2}\,\big\|\nabla\mathcal{L}^{(2)}_{\mathrm{FLC}}\big\|\ \leq\ \big\|\nabla\mathcal{L}^{(2)}_{\mathrm{FLD}}\big\|\ \leq\ L^{2}\,\big\|\nabla\mathcal{L}^{(2)}_{\mathrm{FLC}}\big\|$.
Use $\|J^{\top}y\|\in[m\|y\|,\,L\|y\|]$ (by singular-value bounds) with $y=\mathcal{D}_{a}(\hat{\bm{z}}_{1})-A=\mathcal{D}_{a}(\hat{\bm{z}}_{1})-\mathcal{D}_{a}(\bm{z}_{1})$, and Lemma 1 to get $m\,\|y\|\leq\|J^{\top}y\|\leq L\,\|y\|$ and $m\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|\leq\|y\|\leq L\,\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|$. Multiplying the bounds yields $m^{2}\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|\leq\|J^{\top}y\|\leq L^{2}\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|$. Since $\|\nabla\mathcal{L}^{(2)}_{\mathrm{FLC}}\|=2\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|$, this gives the stated inequality (up to the common factor 2). It follows that step-size sensitivity is governed by the squared condition number $(L/m)^{2}$.
Appendix B Extended Ablations
B.1 Dimensionality Matching for Vision-to-Action Flow
A constraint of flow matching is that the source and target must have identical shapes. In visuomotor contexts, the visual latent representations ($\bm{z}_{0}$) are typically much higher-dimensional than raw action chunks ($A$). A naive solution would be to down-sample the visual representation to match the action dimensionality, which, nevertheless, leads to significant information loss and poor performance, particularly when the dimensional gap is large.
Therefore, we adopt the opposite strategy: we map the raw action chunks into a higher-dimensional latent space that matches the dimensionality of the visual latent representations.
A naive approach is to use fixed linear transformations. When the action dimension is smaller than the latent dimension, we construct a lossless mapping by embedding the actions into the higher-dimensional latent space through zero-padding. The inverse mapping simply discards the padding. Our experiments show that action representations produced by such transformations are insufficient for learning reasonable flow matching policies. We found that learning well-structured action latent spaces via autoencoders as the target distributions for flow matching can be crucial for the success of vision-to-action flow. We develop an action encoder ($\mathcal{E}_{a}$), which does not simply remap dimensions, but also learns structured latent action spaces, making the complex flow from vision to action more tractable.
Table 4: Task SR (%) on ThreadNeedle with different action up-sampling strategies.
Up-sampling Strategy SR (%)
Zero-Padding 0
Action AE (w/o FLD) 0
Action AE (w/ FLD) 92
Autoencoder Loss Selection.
We utilize the L1 loss for the autoencoder loss, $\mathcal{L}_{\text{AE}}$. We found it outperforms the L2 loss, which is prone to mode-averaging and can result in blurry action reconstruction.
Sampling Stochasticity.
We examined VITA variants with different sources of sampling stochasticity. We found that introducing dropout in the network, or variance ($\sigma$) in flow matching (where $\sigma$ injects Gaussian noise into the interpolation path between source and target latents as in OT-CFM) degrades performance. In contrast, adding covariance to the source distribution yields results comparable to the deterministic implementation. Similarly, using a variational objective in the action autoencoder performs on par, whereas applying a variational objective to the image encoder significantly hurts performance (see  Figure˜11).
Appendix C Experiment Results
VITA exhibits efficient and stable learning on AV-ALOHA and PushTFigure˜8) and Robomimic tasks (Figure˜9).
Figure 8: Success rate curves of VITA training on six tasks (five AV-ALOHA + PushT). The curves are mean across three random seeds; the shaded region is $\pm$1 standard deviation.
Figure 9: Success rate curves of VITA training on two Robomimic tasks.
Appendix D Extended Related Works
Imitation Learning for Visuomotor Policy. Imitation learning enables robots to learn complex behaviors by mimicking expert demonstrations. Behavioral cloning, a prominent imitation learning paradigm, frames this as a supervised learning problem, learning a policy that maps observations to actions [26, 27]. Recent advancements in behavioral cloning have widely adopted two modeling methods: generative modeling and autoregressive modeling. Generative methods learn a conditional distribution of actions given an observation. This category includes policies based on conditional variational autoencoders (CVAEs) [26, 27], as well as diffusion [20, 10] and flow matching [30, 14]. On the other hand, autoregressive methods tokenize actions and frame policy learning as a sequence modeling task. These methods predict action tokens sequentially, using next-token prediction [28], next-scale prediction [29], or bi-directional prediction [13]. Generative models ubiquitously require extra conditioning modules (e.g., cross-attention [20], AdaLN [20], FiLM [31, 10]) to inject observations at each step of the generation process. Furthermore, generative and autoregressive methods commonly employ large, expressive networks such as U-Nets or transformers to succeed on complex, high-dimensional robotics tasks. VITA reduces these complexity by formulating the policy as a noise-free and conditioning-free vision-to-action flow; by using vector representations for vision and action, VITA also enables simple and lightweight MLP-only network for policy parameterization, even for challenging bi-manual manipulation tasks.
Diffusion and Flow matching for Generative Modeling. Diffusion [40], grounded in stochastic differential equations (SDEs), generates complex data distributions by sampling from a simple source distribution (typically Gaussian) and iteratively denoising it to the target distribution [19]. Flow matching [18, 41] has been proposed to enable faster training and sampling [41, 32, 42, 18] by modeling the map between source and target distributions with an ordinary differential equation (ODE). Both diffusion and flow matching models have shown strong performance across diverse generative tasks, such as image generation [1, 2, 3, 6, 43, 44], video generation [7, 8], and visuomotor policies [10, 20, 17, 45]. Unlike diffusion, flow matching theoretically places no constraints on the choice of source distribution [32], and a few works have explored leveraging this property to learn the direct transport within the same modality [33, 34], e.g., for image-to-image generation tasks [46, 47]. Recently,  [4] and  [5] extended this to more challenging cross-modal generation between text and image. VITA focuses on learning to bridge vision and action for visuomotor control, where the target modality, action, features sparser data, and lacks semantic structures, compared to text or images, presenting unique challenges. Different from flow matching for image generation, which typically pre-trains and freezes the image encoder and decoder when learning flow matching or diffusion models for image generation [1, 5, 4], VITA resorts to a fully end-to-end pipeline training to effectively learn the latent action space from limited and sparse action data. Furthermore, to enable effective joint training of flow matching and target latent spaces, we propose flow latent decoding to backpropagate action reconstruction losses through the the latent action generation process (ODE solving steps) during training.
Appendix E Discussion of Training Objectives
E.1 Contrastive Latent Alignment
Section˜4.4 introduced two key objectives, FLD and its surrogate FLC, which are both designed to prevent latent space collapse and enable effective VITA learning. Empirically, we find that FLD alone outperforms FLC alone, as it provides a more direct training signal by anchoring the generation process to the ground-truth actions. However, the most robust performance is achieved by combining them, which enforces consistency in both latent space and action space.
Additionally, motivated by the ability of contrastive learning to improve representations and prevent latent collapse [4], we introduced a contrastive loss between vision and action latents [48]. We show that this objective can further boost performance beyond FLD and FLC on some tasks by encouraging the model to learn representations where the similarity between corresponding vision-action pairs is maximized, while the similarity between non-corresponding pairs is minimized.
We employ InfoNCE (Noise-Contrastive Estimation)  [49] for contrastive learning between vision and action. For a given batch of size $N$, the vision latent $\bm{z}_{0,i}$ and action latent $\bm{z}_{1,i}$ from the same data sample are treated as a positive pair. All other non-corresponding combinations $(\bm{z}_{0,i},\bm{z}_{1,j})$ where $i\neq j$ are considered negative pairs. The loss aims to maximize the similarity of positive pairs while minimizing the similarity of negative pairs. The symmetric InfoNCE loss is defined as:
$\mathcal{L}_{\text{contrastive}}=-\frac{1}{2N}\sum_{i=1}^{N}\left[\log\frac{\exp(\text{sim}(\bm{z}_{0,i},\bm{z}_{1,i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(\bm{z}_{0,i},\bm{z}_{1,j})/\tau)}+\log\frac{\exp(\text{sim}(\bm{z}_{1,i},\bm{z}_{0,i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(\bm{z}_{1,i},\bm{z}_{0,j})/\tau)}\right]$
where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity between L2-normalized feature vectors and $\tau$ is a temperature hyperparameter.
As shown in Figure˜10, the contrastive objective, when used alone, is insufficient for effective VITA learning and underperforms VITA with only the FLD loss (see Figure˜6). However, it provides additional performance gains when combined with the FLD and FLC losses, as it helps create a more robust and well-structured latent space.
Figure 10: Success rates using different VITA learning objectives with or without contrastive losses.
E.2 Ablation of the Action Autoencoder
In our experiments, a deterministic autoencoder (AE) learns the target latent space for actions. To investigate the effect of imposing a prior on this latent space, we conducted an ablation study replacing the AE with a variational autoencoder (VAE) [50]. This change introduces a KL divergence regularization term, weighted by $\lambda_{\mathrm{KL}}$, which encourages the encoder’s posterior output, $q(\bm{z}_{1}\!\mid\!A)$, to match a standard normal prior:
$\mathcal{L}_{\mathrm{KL}}\;=\;D_{\mathrm{KL}}\!\big(q(\bm{z}_{1}\mid A)\,\big\|\,p(\bm{z}_{1})\big),\qquad p(\bm{z}_{1})=\mathcal{N}(0,I).$
As shown by the training curves in Figure˜11(a) and  Figure˜12, incorporating this variational objective with various weights ($\lambda_{\mathrm{KL}}$) can degrade model performance.
Figure 11: Ablation of VAE for action (a) and vision (b) on ThreadNeedle.
E.3 Ablation of the Vision Encoder
As described in our main architecture (Figure˜2), the source latent variable $\bm{z}_{0}=\mathcal{E}_{v}(O)\in\mathbb{R}^{D_{\text{latent}}}$ is produced by a deterministic vision encoder. We performed a similar ablation to assess the impact of a prior on this visual latent space. Specifically, we replaced the deterministic encoder with a variational one that models the posterior $q(\bm{z}_{0}\!\mid\!O)$ and introduces a corresponding KL divergence loss (weighted by $\lambda_{\mathrm{KL}}^{\text{obs}}$):
$\mathcal{L}_{\mathrm{KL}}^{\text{obs}}\;=\;D_{\mathrm{KL}}\!\big(q(\bm{z}_{0}\mid O)\,\big\|\,p(\bm{z}_{0})\big),\qquad p(\bm{z}_{0})=\mathcal{N}(0,I).$
As shown in Figure˜11(b), using a VAE for the vision encoder drastically degrades performance.
Appendix F Task Settings
Dataset State Dim Action Dim FPS Image Size Camera
[…]
RoboMimic 43 7 20 $\times256$ agentview_image
Table 5: Comparison of dataset specifications.
F.1 AV-ALOHA Simulation tasks.
CubeTransfer: Pick up a red cube with the right arm and transfer it to the left arm (200 episodes).
SlotInsertion: Use both arms to pick up a green stick and insert it into a pink slot (100 episodes).
HookPackage: Use both arms to pick up a red box and hook it onto a blue wall-mounted hook (100 episodes).
PourTestTube: Pick up two test tubes and pour a small red ball from one into the other (100 episodes).
ThreadNeedle: Pick up a green needle and thread it through the hole of a pink object (200 episodes).
F.2 AV-ALOHA Real tasks
HiddenPick: Lift and open a fabric cover from a box, then pick an object from inside (50 episodes).
TransferFromBox: Pick an object from a box with the right arm, transfer it to the left arm, and place it in another box (50 episodes).
F.3 Robomimic Tasks
Robomimic is a benchmark of single-arm imitation learning tasks [24]. We adapt the environment for compatibility with the LeRobot [38] codebase. The robot state includes arm joint positions (encoded with $\sin$ and $\cos$), joint velocities, end-effector pose, gripper finger positions, and gripper finger velocities. The action space consists of six values for delta position control of the end-effector pose and one value for the absolute position of the gripper.
Square: Pick up a square nut and insert it onto a matching square peg (175 episodes).
Can: Pick up a red can and place it into a box (192 episodes).
Figure 12: Hyperparameter ablations on ThreadNeedle. Each sub-plot varies a single coefficient while holding all others at their default values in  Table˜6. Top: action encoder reconstruction weight, flow latent decoding weight, action encoder contrastive weight. Bottom: latent flow contrastive weight, latent flow consistency weight, action KL weight.
Appendix G Training
In each plot of  Figure˜12, we sweep a single weight while holding all other hyperparameters at the default ( Table˜6), and visualize the success rates over training steps. With these experimental results, a robust configuration on ThreadNeedle includes moderate weights for both AE and FLD (typically in $[0.5,1.0]$), moderate FLC, minimal or no contrastive penalties, and no KL regularization on the action and vision latent.
Appendix H Implementations
H.1 VITA
VITA encodes observations into a latent vector $\bm{z}_{0}$ using a ResNet-18 backbone. The flow network, which learns the mapping from $\bm{z}_{0}$ to $\hat{\bm{z}}_{1}$, is implemented as an MLP. We use a 6-step Euler ODE solver. The latent action $\hat{\bm{z}}_{1}$ is translated to action chunks by a lightweight MLP-based action decoder. A summary of all hyperparameters and loss weights is provided in Table˜6.
H.2 Flow Matching Policy
The FM policy learns a velocity field using a transformer backbone, with AdaLN for conditioning. A 6-step Euler solver is used for solving ODEs. The hyperparameters and loss weights are summarized in Table˜7.
H.3 Diffusion Policy
We follow the DP  [10] implementation in LeRobot [38]. The core of the policy is a U-Net noise predictor, which uses FiLM [31] for conditioning. The model is trained for 100 timesteps using a cosine beta schedule. During inference, trajectories are generated using a 10-step DDPM sampler. All hyperparameters and loss weights are listed in Table˜8.
H.4 Action Chunking Transformer
We follow the ACT  [26] implementation in LeRobot [38]. ACT is a conditional variational autoencoder (cVAE) that generates action chunks conditioned on vision. Its architecture consists of a vision encoder for processing observations and a transformer-based decoder that models the distribution of future actions conditioned on the visual input. A complete list of hyperparameters and loss weights is provided in Table˜9.
Table 6: VITA hyperparameters.
Horizons & Observation
Observation horizon 1
Prediction horizon 16
Action horizon 8
Observer backbone resnet18
Observer tokenize false
VITA core (latent & losses)
Latent dimension ($D_{\text{latent}}$) 512
Decode flow latents true
Consistency weight 1.0
Encoder contrastive weight 1e-4
Flow contrastive weight 0.0
Latent noise std 0.0
Flow matcher / ODE
Flow Matcher OT-CFM
$\sigma$ 0.0
# sampling steps (Euler) 6
Action AE
AE recon loss type l1
Encoder recon weight 0.5
Flow recon (FLD) weight 0.5
Use variational (VAE) false
KL weight (if variational) 1e-6
Freeze encoder / decoder false / false
Pretrained path None
AE network (encoder/decoder)
Encoder type / hidden dim cnn / 512
Decoder type / hidden dim simple / 512
Latent dim (AE) 512
Num heads / MLP ratio 8 / 4
Dropout 0.0
Num layers 4
Table 7: Flow Matching (FM) Policy hyperparameters.
Horizons & Observation
Observation horizon 1
Prediction horizon 16
Action horizon 8
Observer backbone resnet18
Observer tokenize false
Flow matcher / ODE
Flow Matcher Target
$\sigma$ 0.0
# sampling steps 6
Flow network architecture
Backbone flow_transformer
Conditioning adaln  (options: adaln, cross, cross_adaln)
Hidden dim 512
Num layers 4
Num heads 8
MLP ratio 4
Dropout 0.1
Table 8: Diffusion policy hyperparameters.
Horizons & Observation
Observation horizon 1
Prediction horizon 16
Action horizon 8
Observer backbone resnet18
Observer tokenize false
Mask loss for padding false
Diffusion scheduler
Type DDPM
Training timesteps 100
Beta schedule squaredcos_cap_v2
Beta start / end 1e-4 / 2e-2
Prediction type epsilon
Clip sample / range true / 1.0
U-Net architecture
Down dims [512, 1024, 2048] or [256,512,1024]
Kernel size 5
Group norm groups 8
Diffusion step embed dim 128
FiLM scale modulation true
Optimization
Adam LR / backbone scale 1e-4 / 0.1
Adam betas / eps (0.95, 0.999) / 1e-8
Weight decay 1e-6
Scheduler cosine, warmup 500 steps
Training & inference
Total training steps 200000
Inference steps null (defaults to training timesteps)
Table 9: Action Chunking Transformer (ACT): hyperparameters.
Sequence horizons
Action horizon 8
Prediction horizon 16
Observation horizon 1
Observer / backbone
Image encoder ResNet-18
Tokenize false
Transformer (policy head)
Pre-norm false
Model dimension $d_{\text{model}}$ 512
Attention heads 8
Feedforward dim 3200
FFN activation ReLU
Encoder layers 4
Decoder layers 1
Dropout 0.1
Latent / VAE block
Use VAE true
Latent dim 32
VAE encoder layers 4
ACT KL weight 10.0
Optimization
Optimizer Adam
Learning rate $1\times 10^{-5}$
Betas (0.9, 0.999)(0.9,\ 0.999)
$\epsilon$ $1\times 10^{-8}$
Weight decay $1\times 10^{-4}$
Backbone LR scale 0.1
LR scheduler & validation
Scheduler Cosine
Warmup steps 2000
Online validation frequency 2000 steps